Shawniqua is 8 years younger than HiKiem. Their combined ages are 58.

Found 2 solutions by FrankM, Alan3354:
Answer by FrankM(1040)

S= H-8
S+H=58
(H-8)+H=58
H=33
S=25

Answer by Alan3354(69443)

Thx for letting us know.